# Can You Hear Me Now? A Benchmark for Long-Range Graph Propagation

### v1

@@ 1 Introduction
Graphs are fundamental data structures used extensively to represent complex interconnected systems, ranging from social networks and biological pathways, to communication infrastructures and molecular structures. Graph Neural Networks (GNNs) (Sperduti, 1993; Gori et al., 2005; Scarselli et al., 2008; Micheli, 2009; Bruna et al., 2014; Defferrard et al., 2016) have emerged as a successful methodology within deep learning, whose research community was initially driven by the development of diverse architectures capable of capturing intricate relational patterns inherent to graph-structured data, as well as impactful applications across various domains (Hamilton et al., 2017; Derrow-Pinion et al., 2021; Gravina et al., 2022; Gravina & Bacciu, 2024; Khemani et al., 2024).
@@
Currently, we are in the stage in which such pioneer theoretical studies need consolidation, while looking into methodological advancements that can surpass or mitigate such shortcomings. A key enabler of this progress is the establishment of solid and challenging benchmarks that can accurately assess and validate long-range propagation capacities. The availability of controlled synthetic benchmarks, should be complemented by the introduction of compelling application-driven datasets which can clearly demonstrate the practical advantages of addressing long-range propagation issues. Long-range propagation capacities, in this sense, have been noted to be central in key areas of science, such as in biology (Dwivedi et al., 2022; Hariri & Vandergheynst, 2024), biochemistry (Gromiha & Selvaraj, 1999), and climate (Lam et al., 2023).
@@
Existing graph benchmarks have, instead, focused primarily on short to medium-range tasks (Bojchevski & Günnemann, 2018; Shchur et al., 2018; Wu et al., 2018; Sterling & Irwin, 2015; Wale & Karypis, 2006; Hu et al., 2020a; Dwivedi et al., 2023), often overlooking the unique challenges associated with distant information propagation. More recently, the growing interest in this challenge has motivated the community to develop a few benchmarks specifically designed to evaluate information propagation in GNNs. These include the Long-Range Graph Benchmark (LRGB) (Dwivedi et al., 2022) and the Graph Property Prediction (GPP) dataset (Gravina et al., 2023). While this is a significant step forward compared to earlier benchmarks, it does not fully account for the need to capture the true long-range dependencies present in some real-world applications. This is due to limited size of the graphs, the absence of well-defined conditions on the expected propagation range, and the focus of the benchmarks, which is often more aimed at specific issues of over-smoothing and over-squashing, rather than providing a broader evaluation of long-range propagation capabilities. Moreover, LRGB and GPP tasks are facing a natural performance saturation, as novel methodologies are being developed and optimized on them.
@@
Motivated by this, we introduce ECHO (Evaluating Communication over long HOps), a new benchmark designed to assess the capabilities of GNNs to exploit long-range interactions. ECHO consists of three synthetic tasks and two real-world chemically grounded task. The former are designed to provide a controlled setting to assess propagation capabilities. They comprise the prediction of shortest-path-based graph properties (i.e., node eccentricity, single-source shortest paths, and graph diameter) across a diverse graph topologies. These have been defined to increase the difficulty of effective long-range communication, as they present structural bottlenecks for the information flow. The main characteristic of these tasks is that GNNs must heavily rely on global information and effectively learn to traverse the entire graph, similarly to classical algorithms like Bellman-Ford (Bellman, 1958). The real-world tasks target the prediction of molecular total energy and the long-range charge redistribution in molecules, which are critical and practically relevant challenges in computational chemistry (Dupradeau et al., 2010), as they underlie many fundamental processes such as chemical reactivity, molecular stability, and intermolecular interactions. Accurate modeling of these effects is essential for drug design, materials science, and biology understanding.
@@ 2 On the need for a new benchmark
We now elaborate on the need for novel benchmarks specialized on the evaluation of long-range propagation, in relation to existing datasets.
The most widely used benchmark for assessing these capabilities is arguably LRGB (Dwivedi et al., 2022). Its introduction in 2022 has certainly marked an important milestone and promoted the development of the field. However, despite initial rapid improvements, performance on LRGB has now plateaued, showing a noticeable deceleration in progress across the last year, as discussed in Appendix A. In addition to this, it has to be noted that recent works (Tönshoff et al., 2023; Bamberger et al., 2025b) questions the long-range nature of several LRGB tasks, revealing that a subset of tasks is inherently local, rather than requiring long-range diffusion, and that the benchmark itself is highly sensitive to hyperparameter tuning. Other benchmarks propose synthetic tasks on generated structures, including the Tree-Neighborhood (Alon & Yahav, 2021), Graph Property Prediction (Gravina et al., 2023), graph transfer (Di Giovanni et al., 2023; Gravina et al., 2025), GLoRA (Zhou et al., 2025), and Barbell and Clique graphs (Bamberger et al., 2025a). Indeed, most of these tasks are originally designed to address narrow challenges that prevent long-range propagation, such as over-smoothing (Cai & Wang, 2020; Oono & Suzuki, 2020; Rusch et al., 2023) and over-squashing (Alon & Yahav, 2021; Di Giovanni et al., 2023). These phenomena, while related, do not necessarily capture the full spectrum of challenges associated with long-range communication. Moreover, despite being designed to test the ability of GNNs to overcome these limitations, these datasets typically involve small graphs with limited-size diameters. This inherently restricts the propagation radius, creating a significant gap between the benchmark tasks and real-world problems that require much deeper propagation across significantly larger structures.
@@
The ECHO-Charge and ECHO-Energy molecular tasks have strong value per se, proposing a novel, practical, and high-impact challenge for learning models in computational chemistry (Dupradeau et al., 2010). Previous popular benchmarks in this domain (Sterling & Irwin, 2015; Wale & Karypis, 2006; Hu et al., 2020a; Wu et al., 2018; Dwivedi et al., 2022) focused on the prediction of molecular-level properties, such as solubility or HIV inhibition, which are predominantly short-range tasks. This is evident when they can be reduced to the problem of counting small-dimensional local substructures (i.e., with length smaller than 7) (Bouritsas et al., 2023). Differently, ECHO-Charge and ECHO-Energy are the first graph benchmarks that targets long-range interactions at the atomic level, i.e., the microscopic scale. Both benchmarks are not only inherently long-range, but also particularly challenging as they require accurate modeling of charge distributions, energy stabilization, and the complex interplay of atomic interactions. This makes them computationally expensive to solve with current computational chemistry tools. We provide further details on the computational complexity of the underlying quantum simulations in Appendix B.
@@ 3 The ECHO Benchmark
In this section, we introduce a suite of datasets designed to rigorously evaluate the long-range information propagation capabilities of GNNs. Our benchmark consists of two complementary components: a set of algorithmically constructed tasks (collectively called ECHO-Synth) and a set chemically grounded real-world datasets (collectively called ECHO-Chem). Detailed dataset statistics are reported in Appendix D.
@@ 3.1 The ECHO-Synth dataset
These tasks draw inspiration from (Corso et al., 2020; Gravina et al., 2023) [^1] While ECHO-Synth draws inspiration from (Corso et al., 2020; Gravina et al., 2023), it departs from them in a key aspect. Both (Corso et al., 2020) and (Gravina et al., 2023) use small graphs that are mostly sampled from distributions that yield highly connected structures with small diameters and limited long-range dependencies. In contrast, ECHO-Synth is explicitly designed to rigorously stress-test long-range capabilities, leveraging larger graphs and employing topologies deliberately designed to introduce bottlenecks, i.e., requiring substantially more propagation steps for accurate predictions. These design choices significantly increase the long-range difficulty of the tasks, making ECHO-Synth a more rigorous benchmark for evaluating long-range propagation. and were intentionally selected due to their heavy reliance on global information. For example, solving sssp from a given source node requires identifying shortest paths to all other nodes (Dijkstra, 2022), since the information spans the entire graph. Eccentricity builds on this by requiring the longest shortest path from each node, demanding complete graph awareness. Diameter is even more global, involving the longest shortest path between any two nodes (Cormen et al., 2022). Classical algorithms like Dijkstra’s (Dijkstra, 2022) and Bellman-Ford (Bellman, 1958), which perform complete graph traversal, illustrate the challenge these tasks pose for GNNs, which rely on localized message passing. To prevent models from relying on input features rather than learning structural patterns, each node is assigned a uniformly distributed random scalar feature $r\sim\mathcal{U}(0,1)$. [^2] We opted for random uniform node features rather than zero vectors to introduce stochasticity (which makes the synthetic tasks less trivial), and to provide a unique identifier for the nodes in the tasks and prevent the trivial scenario in which all nodes share identical initial representations, hindering the expressiveness of certain architectures (Sato et al., 2021).Additionally, for the sssp task, a binary indicator is included to mark the source node. This ensures that the model can distinguish the source while maintaining uniform input statistics across tasks.
@@
Figure 1: Visualization of the proposed topologies in the synthetic dataset. In all graphs, $N=30$.
@@ 3.2 The ECHO-Chem Datasets
Molecular property prediction is a cornerstone application of GNNs, with common benchmarks involving graph-level prediction tasks such as molecular fingerprint (Duvenaud et al., 2015), solubility, toxicity and various chemical properties (Coley et al., 2017; Hu et al., 2020c). One fundamental task in this domain is the prediction of atomic partial charges, which are continuous, atom-level properties that reflect the electron distribution within a molecule. Accurate charge prediction is essential for modeling molecular interactions, reactivity, and electrostatic behavior. Figure 2 illustrates this task on the 3D molecular graph of caffeine, where each atom is colored according to its predicted partial charge. Complementary to this, another central quantum property of molecular systems is the total energy, which governs stability, chemical reactivity, and conformational preferences. Thus, predicting molecular energies is equally important for chemistry applications.
@@
Traditionally, both atomic charges and molecular energies are computed using quantum mechanical methods, especially Density Functional Theory (DFT) (Argaman & Makov, 2000) or related quantum chemical simulations. While these methods provide high accuracy, their computational cost, arising from solving complex equations, limits their scalability to large molecular datasets or high-throughput tasks. Specifically, high-accuracy simulations require several minutes to process a single molecule. We report a quantitative description of DFT simulation efficiency in Appendix B.
@@
To generate the datasets, we employed a two-step approach. Firstly, the generation process began with molecular 3D structure generation starting from ChEMBL SMILES (Weininger, 1988) strings for all molecules satisfying the given diameter constraint. In order to generate molecular conformations we opted for coordinate optimization using the Generalized Amber Force Field (GAFF) (Grimme et al., 2010), a well-established force field specifically designed for optimizing a wide variety of organic and medically relevant compounds. These optimized structures served as initialization for the subsequent quantum chemical calculations to determine accurate structures, partial charges, and molecular energies. Specifically, we employed Density Functional Theory (DFT) to match the required chemical accuracy required for reliable molecular property annotation. All computations were run with the ORCA package for quantum chemistry (Neese, 2022; Neese et al., 2020; Neese, 2023). A detailed description of the quantum simulations is provided in Appendix B, along with information about the computing platform in Appendix C.
@@ 4 Experiments
Training follows a consistent protocol across all models. We minimize the base-10 logarithm of the Mean Squared Error loss (MSE), $\log_{10}(\mathrm{MSE}(y_{\text{true}}-y_{\text{target}}))$, since the predicted values can be very small in magnitude and this scale-sensitive loss emphasizes small differences. We use the Adam (Kingma & Ba, 2015) optimizer and adopt Early Stopping based on validation loss. with a patience of 100 epochs. The maximum number of training epochs is set to 1000. This procedure ensures convergence while preventing overfitting, and serves as a reference setup to facilitate reproducibility of our results. In order to ensure a fair and robust comparison across all methods and datasets, we employ an extensive hyperparameter optimization protocol. Specifically, for each model-dataset pair, we perform a Bayesian Optimization based on a Gaussian Process prior (Snoek et al., 2012) in the chosen hyperparameter space, spanning 100 trials to explore the respective search space efficiently. We report the complete set of explored hyperparameters for each model, as well as with the selected hyperparameters, in Appendix E. Finally, the best configuration found is validated through four independent training runs, each initialized with a different random seed.
Results on ECHO-Synth dataset. We report results on the synthetic benchmarks in Table 1. All the values are reported using the Mean Absolute Error (MAE). Additional metrics are reported in the Appendix G, Table 18. We start observing that models employing global attention mechanisms significantly outperform traditional message-passing frameworks. Specifically, GRIT demonstrates a superior performance on the sssp task, achieving a remarkably low MAE of $0.121$. In line with literature findings (Dwivedi et al., 2022), this result suggests that incorporating transformer-like global attention substantially mitigates inherent limitations in localized message-passing, which are pronounced in classic architectures such as GCN and GIN. This is further supported by the analysis in Appendix I, which shows that the highest attention scores are often assigned to node pairs that are not directly connected and often far apart in the graph. Interestingly, differential-equation-inspired architectures, particularly those employing non-dissipative or port-Hamiltonian formulations like SWAN, A-DGN, and PH-DGN, consistently perform well across tasks, with similar performance metrics. GRIT achieves the lowest MAE on the diam task ($1.014$), closely followed by SWAN and A-DGN. This highlights the benefit of incorporating attention or non-dissipative dynamics to improve long-range information propagation. Moreover, DRew, reveals its effectiveness in the ecc task, attaining the lowest MAE ($4.651$). This success emphasizes the advantage of multi-hop information propagation, thus effectively addressing topological bottlenecks critical for accurately capturing node eccentricities.
@@
Differently, GraphCON does not inherently outperform traditional methods, and show notably weaker performance relative to other models of the same architectural family (e.g., A-DGN and SWAN). Thus, mere message-passing dynamics that only mitigate the over-smoothing issue does not ensure superior performance in long-range tasks.
@@
Results on ECHO-Chem tasks. We detail the performance of all evaluated models on the atomic partial charge and energy prediction task in Table 2. Additional metrics are reported in the Appendix G, Tables 19 and 20. As anticipated, architectures capable of handling long-range dependencies demonstrate a clear advantage on both benchmarks, given the nature of the task which requires precise modeling of subtle interatomic interactions spread across the molecular graph. Notably, GPS achieves the best performance on ECHO-Energy and it is competitive on ECHO-Charge, confirming the utility of global attention mechanisms in capturing distant influences that modulate quantum chemical properties, albeit at the cost of increased computational complexity (as shown in Appendix H).
Models like A-DGN and SWAN also yield competitive performance, consistently appearing among the top performers, with SWAN emerging as best model in ECHO-Charge. Their success suggests that imposing non-dissipative priors not only improves the propagation dynamics but also guides the model toward chemically plausible solutions. Interestingly, while DRew outperforms classical GNNs, especially in the ECHO-Energy taks, it performs comparatively worse than DE-GNN and transformer-based models.
@@
Traditional message-passing networks, particularly GCN and GIN, again lag behind. These results again confirm the hypothesis that localized aggregation, without mechanisms to to improve propagation effectiveness or integrate distant node information, is inadequate for atomic-level charge modeling. The ECHO-Chem benchmarks thus clearly illustrates the necessity for architectures that either incorporate global attention or embed non-dissipative dynamics to effectively tackle the intricate and non-local dependencies inherent in these molecular tasks.
We provide a visual depiction of charge prediction accuracy on a non-trivial molecule from the test set in Figure 3: the figure compares prediction errors between A-DGN (a) and GCN (b). Atoms are colored by log-error: green = low, orange = high. A-DGN shows consistently lower errors, especially at peripheral atoms, highlighting its ability to capture long-range interactions, while GCN accumulates errors at structurally distant or chemically sensitive sites. This comparison illustrates the benefit of non-dissipative architectures for long-range information propagation on complex graphs and chemical structures.
Lastly, we note that although energies and partial charges errors are small in absolute magnitude across baselines, even subtle deviations (as stated in Dupradeau et al. (2010)), on the order of $10^{-4}\,e$ to $10^{-6}\,e$, can lead to significant downstream effects in molecular modeling and reproducibility of results. Therefore, predictive models must target this level of granularity to produce chemically meaningful outputs.
@@
Additional Experiments and Analysis. We provide further results and a detailed evaluation of baseline performance in Appendix F. Specifically, we investigate how the radius of the explored neighborhood affects each method (Appendix F.1) and how model performance varies across graphs with different diameters (Appendix F.2). Our results consistently indicate that deeper networks outperform shallower ones, confirming the long-range nature of these benchmarks. Moreover, in Appendix F.3 we examine the impact of the readout depth and show that the final performance appears to be independent of this design choice. In Appendix F.4, we further analyze model performance across different graph topologies in ECHO-Synth. The results indicate that, although absolute performance varies slightly with the underlying graph structure, the relative ranking of the models remains consistent, reinforcing the robustness of our findings. In Appendix I, we also visualize GPS’s attention patterns, highlighting the importance of connecting distant nodes to facilitate information flow and improve performance. Together, these analyses reinforce the importance of long-range information propagation in the ECHO benchmark. Finally, Appendix H reports runtime measurements, illustrating a key trade-off between accuracy and efficiency: transformer-based models like GPS achieve strong performance but are computationally demanding, whereas models such as A-DGN provide a more balanced alternative.

### v2

@@ 1 Introduction
Graphs are fundamental data structures used extensively to represent complex interconnected systems, ranging from social networks and biological pathways, to communication infrastructures and molecular structures. Graph Neural Networks (GNNs) (Sperduti, 1993; Gori et al., 2005; Scarselli et al., 2008; Micheli, 2009; Bruna et al., 2014; Defferrard et al., 2016) have emerged as a successful methodology within deep learning, whose research community was initially driven by the development of diverse architectures capable of capturing intricate relational patterns inherent to graph-structured data, as well as impactful applications across various domains (Hamilton et al., 2017; Derrow-Pinion et al., 2021; Gravina et al., 2022; Bacciu et al., 2024; Gravina & Bacciu, 2024; Khemani et al., 2024; Miglior et al., 2025; Tolloso et al., 2026).
@@
Currently, we are in the stage in which such pioneering theoretical studies need consolidation, while looking into methodological advancements that can surpass or mitigate such shortcomings. A key enabler of this progress is the establishment of solid and challenging benchmarks that can accurately assess and validate long-range propagation capacities. The availability of controlled synthetic benchmarks, should be complemented by the introduction of compelling application-driven datasets which can clearly demonstrate the practical advantages of addressing long-range propagation issues. Long-range propagation capacities, in this sense, have been noted to be central in key areas of science, such as in biology (Dwivedi et al., 2022; Hariri & Vandergheynst, 2024), biochemistry (Gromiha & Selvaraj, 1999), and climate (Lam et al., 2023).
@@
Existing graph benchmarks have, instead, focused primarily on short- to medium-range tasks (Bojchevski & Günnemann, 2018; Shchur et al., 2018; Wu et al., 2018; Sterling & Irwin, 2015; Wale & Karypis, 2006; Hu et al., 2020a; Dwivedi et al., 2023), often overlooking the unique challenges associated with distant information propagation. More recently, the growing interest in this challenge has motivated the community to develop a few benchmarks specifically designed to evaluate information propagation in GNNs. These include the Long-Range Graph Benchmark (LRGB) (Dwivedi et al., 2022) and the Graph Property Prediction (GPP) dataset (Gravina et al., 2023). While this is a significant step forward compared to earlier benchmarks, it does not fully account for the need to capture the true long-range dependencies present in some real-world applications. This is due to limited size of the graphs, the absence of well-defined conditions on the expected propagation range, and the focus of the benchmarks, which is often more aimed at specific issues of over-smoothing and over-squashing, rather than providing a broader evaluation of long-range propagation capabilities. Moreover, LRGB and GPP tasks are facing a natural performance saturation, as novel methodologies are being developed and optimized on them.
@@
Motivated by this, we introduce ECHO (Evaluating Communication over long HOps), a new benchmark designed to assess the capabilities of GNNs to exploit long-range interactions. ECHO consists of three synthetic tasks and two real-world chemically grounded tasks. The former are designed to provide a controlled setting to assess propagation capabilities. They comprise the prediction of shortest-path-based graph properties (i.e., node eccentricity, single-source shortest paths, and graph diameter) across diverse graph topologies. These have been defined to increase the difficulty of effective long-range communication, as they present structural bottlenecks to information flow. The main characteristic of these tasks is that GNNs must heavily rely on global information and effectively learn to traverse the entire graph, similarly to classical algorithms like Bellman-Ford (Bellman, 1958). The real-world tasks target the prediction of molecular total energy and the long-range charge redistribution in molecules, which are critical and practically relevant challenges in computational chemistry (Dupradeau et al., 2010), as they underlie many fundamental processes such as chemical reactivity, molecular stability, and intermolecular interactions. Accurate modeling of these effects is essential for drug design, materials science, and biology understanding.
@@ 2 On the need for a new benchmark
We now elaborate on the need for novel benchmarks specialized on the evaluation of long-range propagation in relation to existing datasets.
The most widely used benchmark for assessing these capabilities is arguably LRGB (Dwivedi et al., 2022). Its introduction in 2022 has certainly marked an important milestone and promoted the development of the field. However, despite initial rapid improvements, performance on LRGB has now plateaued, showing a noticeable deceleration in progress across the last year, as discussed in Appendix B. In addition, recent works (Tönshoff et al., 2023; Bamberger et al., 2025b) question the long-range nature of several LRGB tasks, showing that some are inherently local rather than requiring global information, and that the benchmark itself is highly sensitive to hyperparameter tuning. Other benchmarks propose synthetic tasks on generated structures, including the Tree-Neighborhood (Alon & Yahav, 2021), Graph Property Prediction (Gravina et al., 2023), graph transfer (Di Giovanni et al., 2023; Gravina et al., 2025), GLoRA (Zhou et al., 2025), and Barbell and Clique graphs (Bamberger et al., 2025a). Indeed, most of these tasks are originally designed to address narrow challenges that prevent long-range propagation, such as over-smoothing (Cai & Wang, 2020; Oono & Suzuki, 2020; Rusch et al., 2023) and over-squashing (Alon & Yahav, 2021; Di Giovanni et al., 2023). These phenomena, while related, do not necessarily capture the full spectrum of challenges associated with long-range communication. Moreover, despite being designed to test the ability of GNNs to overcome these limitations, these datasets typically involve small graphs with limited-size diameters. This inherently restricts the propagation radius, creating a significant gap between the benchmark tasks and real-world problems that require much deeper propagation across significantly larger structures.
@@
The ECHO-Charge and ECHO-Energy molecular tasks have strong value per se, proposing a novel, practical, and high-impact challenge for learning models in computational chemistry (Dupradeau et al., 2010). Previous popular benchmarks in this domain (Sterling & Irwin, 2015; Wale & Karypis, 2006; Hu et al., 2020a; Wu et al., 2018; Dwivedi et al., 2022) focused on the prediction of molecular-level properties, such as solubility or HIV inhibition, which are predominantly short-range tasks. This is evident when they can be reduced to the problem of counting small-dimensional local substructures (i.e., with length smaller than 7) (Bouritsas et al., 2023). Differently, ECHO-Charge and ECHO-Energy are the first graph benchmarks that target long-range interactions at the atomic level, i.e., the microscopic scale. Both benchmarks are not only inherently long-range but also particularly challenging, as they require accurate modeling of charge distributions, energy stabilization, and the complex interplay of atomic interactions. This makes them computationally expensive to solve with current computational chemistry tools. We provide further details on the computational complexity of the underlying quantum simulations in Appendix C.
@@ 3 The ECHO Benchmark
In this section, we introduce a suite of datasets designed to rigorously evaluate the long-range information propagation capabilities of GNNs. Our benchmark consists of two complementary components: a set of algorithmically constructed tasks (collectively called ECHO-Synth) and a set chemically grounded real-world datasets (collectively called ECHO-Chem). Detailed dataset statistics are reported in Appendix E.
@@ 3.1 The ECHO-Synth dataset
These tasks draw inspiration from Corso et al. (2020); Gravina et al. (2023) [^1] While ECHO-Synth draws inspiration from Corso et al. (2020); Gravina et al. (2023), it departs from them in a key aspect. Both Corso et al. (2020) and Gravina et al. (2023) use small graphs that are mostly sampled from distributions that yield highly connected structures with small diameters and limited long-range dependencies. In contrast, ECHO-Synth is explicitly designed to rigorously stress-test long-range capabilities, leveraging larger graphs and employing topologies deliberately designed to introduce bottlenecks, i.e., requiring substantially more propagation steps for accurate predictions. These design choices significantly increase the long-range difficulty of the tasks, making ECHO-Synth a more rigorous benchmark for evaluating long-range propagation. and were intentionally selected due to their heavy reliance on global information. For example, solving sssp from a given source node requires identifying shortest paths to all other nodes (Dijkstra, 2022), since the information spans the entire graph. Eccentricity builds on this by requiring the longest shortest path from each node, demanding complete graph awareness. Diameter is even more global, involving the longest shortest path between any two nodes (Cormen et al., 2009). Classical algorithms like Dijkstra’s (Dijkstra, 2022) and Bellman-Ford (Bellman, 1958), which perform complete graph traversal, illustrate the challenge these tasks pose for GNNs, which rely on localized message passing. To prevent models from relying on input features rather than learning structural patterns, each node is assigned a uniformly distributed random scalar feature $r\sim\mathcal{U}(0,1)$. [^2] We opted for random uniform node features rather than zero vectors to introduce stochasticity (which makes the synthetic tasks less trivial), and to provide a unique identifier for the nodes in the tasks and prevent the trivial scenario in which all nodes share identical initial representations, hindering the expressiveness of certain architectures (Sato et al., 2021).Additionally, for the sssp task, a binary indicator is included to mark the source node. This ensures that the model can distinguish the source while maintaining uniform input statistics across tasks.
@@
Figure 1: Visualization of the proposed topologies in the synthetic dataset. All graphs consist of 30 nodes to ease visualization, and different node colors correspond to different random initial features.
@@ 3.2 The ECHO-Chem Datasets
Molecular property prediction is a cornerstone application of GNNs, with common benchmarks involving graph-level prediction tasks such as molecular fingerprint (Duvenaud et al., 2015), solubility, toxicity and various chemical properties (Coley et al., 2017; Hu et al., 2020b). One fundamental task in this domain is the prediction of atomic partial charges, which are continuous, atom-level properties that reflect the electron distribution within a molecule. Accurate charge prediction is essential for modeling molecular interactions, reactivity, and electrostatic behavior. Figure 2 illustrates this task on the 3D molecular graph of caffeine, where each atom is colored according to its predicted partial charge. Complementary to this, another central quantum property of molecular systems is the total energy, which governs stability, chemical reactivity, and conformational preferences. Thus, predicting molecular energies is equally important for chemistry applications.
@@
Traditionally, both atomic charges and molecular energies are computed using quantum mechanical methods, especially Density Functional Theory (DFT) (Argaman & Makov, 2000) or related quantum chemical simulations. While these methods provide high accuracy, their computational cost, arising from solving complex equations, limits their scalability to large molecular datasets or high-throughput tasks. Specifically, high-accuracy simulations require several minutes to process a single molecule. We report a quantitative description of DFT simulation efficiency in Appendix C.
@@
To generate the datasets, we employed a two-step approach. Firstly, the generation process began with molecular 3D structure generation starting from ChEMBL SMILES (Weininger, 1988) strings for all molecules satisfying the given diameter constraint. In order to generate molecular conformations we opted for coordinate optimization using the Generalized Amber Force Field (GAFF) (Grimme et al., 2010), a well-established force field specifically designed for optimizing a wide variety of organic and medically relevant compounds. These optimized structures served as initialization for the subsequent quantum chemical calculations to determine accurate structures, partial charges, and molecular energies. Specifically, we employed Density Functional Theory (DFT) to match the required chemical accuracy required for reliable molecular property annotation. All computations were run with the ORCA package for quantum chemistry (Neese, 2022; Neese et al., 2020; Neese, 2023). A detailed description of the quantum simulations is provided in Appendix C, along with information about the computing platform in Appendix D.
@@ 4 Experiments
Training follows a consistent protocol across all models. We minimize the base-10 logarithm of the Mean Squared Error loss (MSE), $\log_{10}(\mathrm{MSE}(y_{\text{true}}-y_{\text{target}}))$, since the predicted values can be very small in magnitude and this scale-sensitive loss emphasizes small differences. We use the Adam (Kingma & Ba, 2015) optimizer and adopt Early Stopping based on validation loss with a patience of 50 epochs. The maximum number of training epochs is set to 1000. This procedure ensures convergence while preventing overfitting, and serves as a reference setup to facilitate reproducibility of our results. In order to ensure a fair and robust comparison across all methods and datasets, we employ an extensive hyperparameter optimization protocol. Specifically, for each model-dataset pair, we perform a Bayesian Optimization based on a Gaussian Process prior (Snoek et al., 2012) in the chosen hyperparameter space, spanning 100 trials to explore the respective search space efficiently. We report the complete set of explored hyperparameters for each model, as well as with the selected hyperparameters, in Appendix F. Finally, the best configuration found is validated through four independent training runs, each initialized with a different random seed.
Results on ECHO-Synth dataset.
@@
We report results on the synthetic benchmarks in Table 1. All the values are reported using the Mean Absolute Error (MAE). Additional metrics are reported in the Appendix H, Table 18. We start observing that models employing global attention mechanisms significantly outperform traditional message-passing frameworks. Specifically, GRIT demonstrates a superior performance on the sssp task, achieving a remarkably low MAE of $0.121$. In line with literature findings (Dwivedi et al., 2022), this result suggests that incorporating transformer-like global attention substantially mitigates inherent limitations in localized message-passing, which are pronounced in classic architectures such as GCN and GIN. This is further supported by the analysis in Appendix J, which shows that the highest attention scores are often assigned to node pairs that are not directly connected and often far apart in the graph. Interestingly, differential-equation-inspired architectures, particularly those employing non-dissipative or port-Hamiltonian formulations (SWAN, A-DGN, and PH-DGN), consistently perform well across tasks, with similar performance metrics. GRIT achieves the lowest MAE on the diam task ($1.014$), closely followed by SWAN and A-DGN. This highlights the benefit of incorporating attention or non-dissipative dynamics to improve long-range information propagation. Moreover, DRew, reveals its effectiveness in the ecc task, attaining the lowest MAE ($4.651$). This success emphasizes the advantage of multi-hop information propagation, thus effectively addressing topological bottlenecks critical for accurately capturing node eccentricities. Differently, GraphCON does not inherently outperform traditional methods, and show notably weaker performance relative to other models of the same architectural family (e.g., A-DGN and SWAN). Thus, mere message-passing dynamics that only mitigate the over-smoothing issue does not ensure superior performance in long-range tasks.
@@
Results on ECHO-Chem tasks.
@@
We detail the performance of all evaluated models on the atomic partial charge and energy prediction task in Table 2. Additional metrics are reported in the Appendix H, Tables 19 and 20. As anticipated, architectures capable of handling long-range dependencies demonstrate a clear advantage on both benchmarks, given the nature of the task which requires precise modeling of subtle interatomic interactions spread across the molecular graph.
Notably, GPS achieves the best performance on ECHO-Energy and it is competitive on ECHO-Charge, confirming the utility of global attention mechanisms in capturing distant influences that modulate quantum chemical properties, albeit at the cost of increased computational complexity (as shown in Appendix I).
Models like A-DGN and SWAN also produce competitive performance, appearing consistently among the top performers, with SWAN emerging as the best model in ECHO-Charge. Their success suggests that imposing non-dissipative priors not only improves the propagation dynamics but also guides the model toward chemically plausible solutions. Interestingly, while DRew outperforms classical GNNs, especially in the ECHO-Energy taks, it performs comparatively worse than DE-GNN and transformer-based models.
Traditional message-passing networks, particularly GCN and GIN, again lag behind. These results again confirm the hypothesis that localized aggregation, without mechanisms to improve propagation effectiveness or integrate distant node information, is inadequate for atomic-level charge modeling. The ECHO-Chem benchmarks thus clearly illustrate the necessity for architectures that either incorporate global attention or embed non-dissipative dynamics to effectively tackle the intricate and non-local dependencies inherent in these molecular tasks.
We provide a visual depiction of charge prediction accuracy on a non-trivial molecule from the test set in Figure 3: the figure compares prediction errors between GPS and GCN. GPS shows consistently lower errors, especially at peripheral atoms, highlighting its ability to capture long-range interactions, while GCN accumulates errors at structurally distant or chemically sensitive sites. This comparison illustrates the benefit of long-range information propagation on complex graphs and chemical structures.
Lastly, we note that although the energies and partial charges errors are small in absolute magnitude across baselines, even subtle deviations (as stated in Dupradeau et al. (2010)), on the order of $10^{-4}\,e$ to $10^{-6}\,e$, can lead to significant downstream effects in molecular modeling and reproducibility of results. Therefore, predictive models must target this level of granularity to produce chemically meaningful output.
@@
Additional Experiments and Analysis. We provide further results and a detailed evaluation of baseline performance in Appendix G. Specifically, we investigate how the radius of the explored neighborhood affects each method (Appendix G.1) and how model performance varies across graphs with different diameters (Appendix G.2). Our results consistently indicate that deeper networks outperform shallower ones, confirming the long-range nature of these benchmarks. Moreover, in Appendix G.3 we examine the impact of the readout depth and show that the final performance appears to be independent of this design choice. In Appendix G.4, we further analyze model performance across different graph topologies in ECHO-Synth. The results indicate that, although absolute performance varies slightly with the underlying graph structure, the relative ranking of the models remains consistent, reinforcing the robustness of our findings. In Appendix J, we also visualize GPS’s attention patterns, highlighting the importance of connecting distant nodes to facilitate information flow and improve performance. Together, these analyses reinforce the importance of long-range information propagation in the ECHO benchmark. Finally, Appendix I reports runtime measurements, illustrating a key trade-off between accuracy and efficiency: transformer-based models like GPS achieve strong performance but are computationally demanding, whereas models such as A-DGN provide a more balanced alternative.